OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (–1, 8)?

OpenStudy (anonymous):

\[M=\frac{y1-y2}{x1-x2}=\frac{5-8}{2--1}=-1\]

OpenStudy (anonymous):

y=Mx+b...........y=-1x+b.............5=-1(2)+b.........(plugged in 2,5 for x and y)

OpenStudy (anonymous):

its simple

OpenStudy (anonymous):

b=7

OpenStudy (anonymous):

the equation y=-1x+7

OpenStudy (anonymous):

none of those of are my answers

OpenStudy (anonymous):

A. y – 8 = 1(x – 2) B. y + 8 = –1(x – 1) C. y + 5 = 1(x + 2) D. y – 5 = –1(x – 2)

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

Given a point \[(a,b)\]the point slope form is\[y-b=m(x-a)\]Where m is the slope. Recall that slope is rise over run, so for two arbitrary points\[(x_1, y_1) \text{ and } (x_2,y_2)\]the slope, rise over run, is given by\[m=\dfrac{y_2-y_1}{x_2-x_1}\]You can use either of these points as (a,b) above.

OpenStudy (anonymous):

\[(y-5)/(x-2)=(5-8)/(2+1)\] (y-2)/(x-2)=-1 (y-5)=-1(x-2) ans D